tara is eight years younger than her sister anna.
if the sum of their age is 34 how old is tara

Answer by lenny460(1073)

Let their ages be:
Anna = n
Tara = n - 8
Therefore:
n + (n - 8) = 34
n + n - 8 = 34
Combine like terms:
n + n = 2n
2n - 8 = 34
Add 8 to each side of the equation.
2n - 8 + 8 = 34 + 8
Combine like terms:
-8 + 8 = 0
34 + 8 = 42
2n = 42
Divide each side by 2
2n/2 = 42/2
n = 21
n - 8 = 21 - 8 = 13
Anna is 21 years old and
Tara is 13 years old.


obuong3@aol.com